# Enhancing Adversarial Transferability via Component-Wise Augmentation Method

### v1

@@ Abstract
Deep Neural Networks (DNNs) are highly vulnerable to adversarial examples, which pose significant challenges in security-sensitive applications. Among various adversarial attack strategies, input transformation-based attacks have demonstrated remarkable effectiveness in enhancing adversarial transferability. However, existing methods fail to diversify attention regions across models adequately and introduce excessive information loss during transformations. In this paper, we introduce a novel input transformation-based method, termed Component-Wise Augmentation (CWA), designed to enhance transferability by locally applying block-wise transformations. CWA strategically integrates interpolation and selective rotation on individual image blocks to diversify model attention regions while preserving semantic integrity. Extensive experiments on the standard ImageNet dataset show that CWA consistently outperforms state-of-the-art methods in both attack success rates and stability across CNN- and Transformer-based models, while also demonstrating superior performance against multiple defense methods.
@@ 1 Introduction
Figure 1: The heatmaps generated on ResNet-18 of raw image and its transformed images by DIM, SIM, Adimx, MaskBlock, US-MM, BSR, and our proposed CWA.
@@
To improve adversarial transferability, researchers have proposed a variety of methods for generating adversarial examples. Gradient-based methods [9, 16, 4, 20, 8] leverage the gradient information of models to craft adversarial examples with high computational efficiency and serve as the foundation for many other approaches. However, they generally suffer from poor transferability, particularly across models with different architectures. Model-related methods [47, 46, 24] exploit internal model features, such as skip connections or backpropagation, to generate architecture-specific adversarial examples. While these methods perform well for specific architectures, they often overfit and lack generalizability, making them less effective for diverse architectures. Ensemble-based [21, 51, 17, 1] and generation-based [33] methods aggregate the outputs of multiple models or use generative models like GANs to directly generate adversarial examples. While these approaches can improve transferability, they are constrained by high computational costs, reliance on joint training of multiple models, and limited scalability. In contrast, input transformation-based methods [49, 49, 20, 53] enhance cross-model transferability through enriching adversarial example diversity with operations such as cropping, scaling, and rotation. These methods are computationally efficient and do not require access to model parameters, making them particularly suitable for black-box attack scenarios.
However, current input transformation-based methods still face critical challenges. As shown in Figure 1, some methods fail to effectively shift the model’s attention, while others redirect attention to irrelevant regions of the object, introducing significant information loss in the process.
To address these issues, we propose a novel method called Component-Wise Augmentation (CWA), which enhances the transferability of adversarial examples by encouraging models to focus on diverse regions of the object in the original image. Specifically, CWA applies block-wise transformations that interpolate and selectively rotate image patches while preserving essential semantic information. By introducing these localized transformations, our method generates adversarial examples with enriched attention distributions that generalize effectively across different model architectures.
@@
    A New Perspective: We provide a fresh perspective on solving the untargeted transferability challenge. Instead of broadly enriching input images, our approach focuses on crafting transformed images that maximize the diversity of attention regions within the object area in original input images.
@@
    Novel Input Transformation Framework: We propose CWA, an innovative method combining interpolation and selective rotation to enhance adversarial transferability by preserving semantic integrity while diversifying attention distributions.
@@
    State-of-the-Art Performance: Our method achieves state-of-the-art results on the ImageNet dataset, demonstrating superior attack success rates and lower standard deviations compared to existing approaches.
@@ 2.1 Input Transformation-based Attacks
Input transformation-based methods have been widely studied to enhance adversarial transferability, particularly in black-box settings. These methods diversify the input samples used for gradient computation, enabling adversarial examples to generalize across different model architectures. We classify these methods into four categories: scale-based methods, mixed image-based methods, block-based methods, and adaptive-based methods. Scale-based methods [49, 20] focus on resizing and transforming input images to introduce diversity in gradient computation. Mixed image-based methods [41, 39, 53] enrich input diversity by incorporating features from other images or categories. However, whole-image transformations alone may not effectively alter the model’s attention. To address this, Block-based methods [7, 38] focus on partitioning input images into blocks and applying transformations to each block independently. Furthermore, to better utilize existing transformation methods, adaptive-based methods [45, 43, 54] introduce learnable transformation strategies to optimize adversarial transferability.
@@ 3 Methodology
In this section, we first introduce the prerequisite knowledge and our motivation. Then, we provide a detailed explanation of our CWA method. To highlight the differences between our approach and other methods, we also summarize the distinctions between our method and DIM [49], as well as BSR [38].
@@ 3.2 Motivation
3.3 Component-Wise Augmentation
@@
  • Maximize the retention of original image information: Ensure that essential features and semantics of the original image are preserved to maintain the model’s performance.
@@
To overcome these limitations, we propose CWA, which applies a sequence of transformations to individual blocks of an image. These transformations include interpolation (for more details on interpolation, refer to Appendix 7), comprising pre-interpolation (interpolation-based shrinking) and block-wise scaling (interpolation-based enlargement), as well as selective rotation. Together, these operations enhance local feature diversity and improve the model’s robustness against adversarial perturbations.
Formally, the CWA process involves the following steps:
@@
    Selective Rotation: To further diversify the spatial representation of features and avoid the massive information loss caused by rotating all blocks, Selective Rotation is applied to a subset of blocks. Let $N$ be the total number of blocks, and let $\mathcal{R}\subseteq\{1,2,\dots,N\}$ denote the indices of $k$ randomly selected blocks for rotation ($|\mathcal{R}|=k$).
    For each selected block $B_{i,j}\in\mathcal{R}$, a random rotation angle $r_{i,j}$ is sampled from:
@@
    $T_{k}(I)=\bigcup_{i=1}^{n}\bigcup_{j=1}^{n}\tilde{B}_{i,j}.$ (10)
To stabilize gradient calculations during adversarial training with CWA, we compute the average gradient over $N$ transformed images:
@@
Where $T(x,n,N,s_{\text{max}},k,r)$ represents the transformation applied to the augmented image, parameterized by the input image $x$, the number of blocks $n$, the total number of transformed images $N$, the maximum scaling factor $s_{\text{max}}$ (with $s_{\text{min}}=1.0$), the maximum rotation angle $r$, and the number of rotated blocks $k$. $J$ denotes the loss function (e.g., cross-entropy), and $y$ is the ground truth label.
Here we integrate our CWA method into MI-FGSM, and summarize the algorithm in Algorithm 1
Algorithm 1 Component-Wise Augmentation
@@
6:     Clip $\delta$ to ensure $\|\delta\|_{\infty}\leq\epsilon$
7:     Update the adversarial example by:
@@
3.4 Comparison of CWA with DIM and BSR
DIM enhances the transferability of adversarial attacks by applying resizing and padding to input samples, whereas BSR achieves this by randomly shuffling and rotating image blocks. Despite incorporating similar operations, our method, CWA, differs significantly from these approaches. Below, we highlight the distinctions between our method and the two techniques.
  • CWA Vs. DIM In contrast to DIM, which performs scaling with fixed values on the entire image, our method emphasizes a random enlargement operation with pre-interpolation. Moreover, while DIM applies this operation to the entire image uniformly, our approach targets individual blocks within the image, introducing more localized transformations.
  • CWA Vs. BSR As for BSR, although both methods involve block partitioning, our approach discards the shuffle operation and primarily utilizes block interpolation to modify the image structure. Regarding rotation, instead of applying rotation to all blocks, our method adopts a selective rotation mechanism, where the number of rotated blocks is limited.
@@ 4 Experiments
Model Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin Mean Std. Dev.
ResNet-18 DIM 100* 61.7 66.1 90.4 30.4 37.4 53.4 56.9 62.0 23.8
@@ 4.1 Experimental Setup
Models. We evaluate our proposed CWA method across three categories of target models: 1) CNN-based models, including four widely recognized architectures: ResNet-18 [11], ResNet-101 [11], ResNeXt-50 [50], and DenseNet-121 [14]; 2) Transformer-based models, comprising ViT-B [5], PiT [12], Visformer [2], and Swin [22]; 3) Defense models, which include five defense methods: AT [35], HGD [19], neural NRP [26], RS [3], and DiffPure [27]; All models are pre-trained on the ImageNet dataset and evaluated on single model.
Dataset. Following previous works [^1] https://github.com/Trustworthy-AI-Group/TransferAttack, We evaluate our proposed CWA on 1000 images belonging to 1000 categories from the validation set of ImageNet dataset. All images are classified correctly by the models.
Baselines. We compare CWA with other input transformation-based methods. Specifically, the image scale-based methods (DIM [49], SIM [20]), the mixed image-based methods (Admix [41], US-MM [39]), and the block-based methods (MaskBlock [7], BSR [38]). For fairness, all the input transformations are integrated into MI-FGSM [4].
Evaluation Settings. We set the maximum perturbation $\epsilon=16/255$, the number of iterations $epoch=10$, the step size $\alpha=\epsilon/epoch$, the batch size $batchsize=8$, and the decay factor $\mu=1$ for MI-FGSM. For our method, CWA generates 20 scaled copies per iteration, divides the image into 2x2 blocks, applies a scaling factor ranging from 1.0 to 1.3, and applies a maximum rotation angle of $26^{\circ}$, selectively rotating $k=2$ blocks. For other methods, we follow the parameters reported in the original papers. But for US-MM, we modify the number of scaled copies to 20 per iteration for fairness.
@@ 4.2 Evaluations on CNN-based Models
We first evaluate the transferability of adversarial examples generated by various attacks with input transformations when the surrogate model is based on a CNN architecture. Specifically, we generate adversarial examples using ResNet-18, ResNet-101, ResNeXt-50, and DenseNet-121 as surrogate models and evaluate their attack success rates (ASR) across eight target models. The results are summarized in Table 1, where the first column indicates the surrogate model, the second column lists the attack methods, and the remaining columns present the ASR of different classification models under attack and the last two columns are the mean ASR and standard deviation, respectively.
Our proposed method consistently achieves state-of-the-art performance, with the highest mean ASR and lowest standard deviation across almost all experimental settings. Notably, when the surrogate model is ResNet-18, our method achieves the highest ASR across all target models. Furthermore, when the target models are ViT and PiT, our method surpasses the previous state-of-the-art BSR more than 6.5% and outperforms other baseline methods by at least 24%.
These results confirm the effectiveness of our approach in generating transferable adversarial examples when the surrogate model is CNN-based.
@@ 4.3 Evaluations on Transformer-based models
Model Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin Mean Std. Dev.
ViT DIM 56.9 47.8 49.6 59.7 89.5* 54.0 55.0 61.7 59.3 13.1
@@
Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin
@@
Attack ResNet-18 ResNet-101 ResNeXt-50 DenseNet-121 ViT PiT Visformer Swin
@@
CWA 40.5 56.7 38.4 41.3 43.2 43.2 48.7 37.8
Table 4: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under Diffure. The classifier is ResNet-101.
Building on the results observed with CNN-based models, we next evaluate our method’s performance when the surrogate model is based on a transformer architecture. Specifically, we use transformer-based models, including ViT, PiT, Visformer, and Swin, as surrogate models to generate adversarial examples and evaluate their transferability. The results are summarized in Table 2.
From the results, we observe that our method also achieves SOTA performance in terms of both mean ASR and standard deviation across most of experimental settings when the surrogate model is transformer-based. Notably, when the surrogate model is ViT, our method does not achieve the highest ASR on the original classification model (ViT itself) but consistently achieves the best performance on all other classification models. This demonstrates the robustness and generalizability of our approach, as it effectively transfers adversarial examples to different target architectures, even when the surrogate model has inherent challenges in transferability.

### v2

# Enhancing Adversarial Transferability via Component-Wise Transformation

@@ Abstract
Deep Neural Networks (DNNs) are highly vulnerable to adversarial examples, which pose significant challenges in security-sensitive applications. Among various adversarial attack strategies, input transformation-based attacks have demonstrated remarkable effectiveness in enhancing adversarial transferability. However, existing methods still perform poorly across different architectures, even though they have achieved promising results within the same architecture. This limitation arises because, while models of the same architecture may focus on different regions of the object, the variation is even more pronounced across different architectures. Unfortunately, current approaches fail to effectively guide models to attend to these diverse regions. To address this issue, this paper proposes a novel input transformation-based attack method, termed Component-Wise Transformation (CWT). CWT applies interpolation and selective rotation to individual image blocks, ensuring that each transformed image highlights different target regions, thereby improving the transferability of adversarial examples. Extensive experiments on the standard ImageNet dataset show that CWT consistently outperforms state-of-the-art methods in both attack success rates and stability across CNN- and Transformer-based models.
@@ 1 Introduction
Figure 1: There shows our CWT process for image transformation. The heatmaps generated on ResNet-18. Our method successfully enables a single surrogate model to focus on different regions of an object.
@@
To improve adversarial transferability, researchers have proposed a variety of methods for generating adversarial examples. Gradient-based methods [10, 17, 4, 20, 8] leverage the gradient information of models to craft adversarial examples with high computational efficiency and serve as the foundation for many other approaches. However, they generally suffer from poor transferability, particularly across models with different architectures. Model-related methods [48, 47, 24] exploit internal model features, such as skip connections or backpropagation, to generate architecture-specific adversarial examples. While these methods perform well for specific architectures, they often overfit and lack generalizability. Ensemble-based [21, 52, 18, 1] and generation-based [33] methods aggregate the outputs of multiple models or use generative models like GANs to directly generate adversarial examples. While these approaches can improve transferability, they are constrained by high computational costs. In contrast, input transformation-based methods [50, 50, 20, 54] enhance cross-model transferability through enriching adversarial example diversity with operations such as cropping, scaling, and rotation. These methods are computationally efficient and do not require access to model parameters, making them particularly suitable for black-box attack scenarios.
Figure 2: The heatmaps generated on ResNet-18 of raw image and its transformed images by DIM, SIM, Adimx, MaskBlock, US-MM, BSR, and our proposed CWT. Notably, CWT successfully shifts the surrogate model’s attention toward the central region of the object.
However, existing input transformation-based methods still face critical challenges. As illustrated in Figure 2, some approaches fail to effectively alter the model’s attention, resulting in an attention distribution that remains consistent with the original image. In contrast, others misdirect attention to irrelevant regions of the object, deviating from the areas of interest in the original image.
To address these issues, we propose a novel method called Component-Wise Transformation (CWT), which enhances the transferability of adversarial examples by encouraging models to focus on diverse regions of the object in the original image. Specifically, CWT applies block-wise transformations that interpolate and selectively rotate image patches while preserving essential semantic information. By introducing these localized transformations, our method generates adversarial examples with enriched attention distributions that generalize effectively across different model architectures.
@@
    We introduce a novel perspective on addressing the challenge of untargeted adversarial transferability. Instead of broadly enriching input images, our approach focuses on generating transformed images that enable a single surrogate model to attend to different regions of the object across various transformed images.
@@
    We propose CWT, an innovative method combining interpolation and selective rotation to enhance adversarial transferability by diversifying attention distributions.
@@
    Our method achieves state-of-the-art results on the ImageNet dataset, demonstrating superior attack success rates and lower standard deviations compared to existing approaches.
@@ 2.1 Input Transformation-based Attacks
Input transformation-based methods have been widely studied to enhance adversarial transferability, particularly in black-box settings. These methods diversify the input samples used for gradient computation, enabling adversarial examples to generalize across different model architectures. We classify these methods into four categories: scale-based methods, mixed image-based methods, block-based methods, and adaptive-based methods. Scale-based methods [50, 20] focus on resizing and transforming input images to introduce diversity in gradient computation. Mixed image-based methods [42, 40, 54] enrich input diversity by incorporating features from other images. However, whole-image transformations alone may not effectively alter the model’s attention. To address this, Block-based methods [7, 39] focus on partitioning input images into blocks and applying transformations to each block independently. Furthermore, to better utilize existing transformation methods, adaptive-based methods [46, 44, 56] introduce learnable transformation strategies to optimize adversarial transferability.
@@ 3 Methodology
In this section, we first introduce the prerequisite knowledge and our motivation. Then, we provide a detailed explanation of our CWT method. To highlight the differences between our approach and other methods, we also summarize the distinctions between our method and DIM [50], as well as BSR [39].
@@ 3.2 Motivation
3.3 Component-Wise Transformation
@@
  • Maximize the retention of original image information: Ensure that essential features and semantics of the original image are preserved.
@@
To overcome these limitations, we propose CWT, which applies a sequence of transformations to individual blocks of an image. As shown in Figure 1, these transformations include bilinear interpolation (for more details on interpolation, refer to Appendix B), comprising pre-interpolation (interpolation-based shrinking) and block-wise scaling (interpolation-based enlargement), as well as selective rotation.
Formally, the CWT process involves the following steps:
@@
    Selective Rotation: To further diversify the spatial representation of features and avoid the massive information loss caused by rotating all blocks, Selective Rotation is applied to a subset of blocks. Let $N$ be the total number of blocks, and let $\mathcal{R}\subseteq\{1,2,\dots,N\}$ denote the indices of $k$ randomly selected blocks for rotation ($|\mathcal{R}|=k$). For each selected block $B_{i,j}\in\mathcal{R}$, a random rotation angle $r_{i,j}$ is sampled from:
@@
    $R_{k}(I)=\bigcup_{i=1}^{n}\bigcup_{j=1}^{n}\tilde{B}_{i,j}.$ (10)
To stabilize gradient calculations during adversarial training with CWT, we compute the average gradient over $N$ transformed images:
@@
Where $T(x^{\text{adv}},n,N,s_{\text{max}},k,r)$ represents the transformation applied to the augmented image, parameterized by the input image $x^{\text{adv}}$, the number of blocks $n$, the total number of transformed images $N$, the maximum scaling factor $s_{\text{max}}$ (with $s_{\text{min}}=1.0$), the maximum rotation angle $r$, and the number of rotated blocks $k$. $J$ denotes the loss function (cross-entropy), and $y$ is the ground truth label.
Here we integrate our CWT method into MI-FGSM, and summarize the algorithm in Algorithm 1.
Algorithm 1 Component-Wise Transformation
@@
6:     Update the adversarial example by:
@@
7:     Clip $\delta$ to ensure $\|\delta\|_{\infty}\leq\epsilon$:
$x^{\text{adv}}_{t}=\text{Clip}_{x,\epsilon}(x^{\text{adv}}_{t}-x,-\epsilon,\epsilon)$
@@
3.4 Comparison of CWT with DIM and BSR
DIM enhances the transferability of adversarial attacks by applying resizing and padding to input samples, whereas BSR achieves this by randomly shuffling and rotating image blocks. Despite incorporating similar operations, our method, CWT, differs significantly from these approaches. Below, we highlight the distinctions between our method and the two methods.
  • CWT Vs. DIM In contrast to DIM, which performs scaling with fixed values on the entire image, our method emphasizes a random enlargement operation with pre-interpolation. Moreover, while DIM applies this operation to the entire image uniformly, our approach targets individual blocks within the image, introducing more localized transformations.
  • CWT Vs. BSR As for BSR, although both methods involve block partitioning, our approach discards the shuffle operation and primarily utilizes block interpolation to modify the image structure. Regarding rotation, instead of applying rotation to all blocks, our method adopts a selective rotation mechanism, where the number of rotated blocks is limited.
@@ 4 Experiments
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
RN-18 DIM 100* 61.7 66.1 90.4 30.4 37.4 53.4 56.9 62.0 23.8
@@ 4.1 Experimental Setup
Models. We evaluate our proposed CWT method across three categories of target models: 1) CNN-based models, including four widely recognized architectures: ResNet-18 (RN-18) [12], ResNet-101 (RN-101) [12], ResNeXt-50 (RX-50) [51], and DenseNet-121 (DN-121) [15]; 2) Transformer-based models, comprising ViT-B [5], PiT-B [13], Visformer (ViF-S) [2], and Swin-Tiny (Swin-T) [22]; 3) Defense models, which include four defense methods: AT [35], HGD [19], RS [3], and DiffPure [27]; All models are pre-trained on the ImageNet dataset and evaluated on single model.
Dataset. Following previous works [^1] https://github.com/Trustworthy-AI-Group/TransferAttack, We evaluate our proposed CWT on 1000 images belonging to 1000 categories from the validation set of ImageNet dataset. All images are classified correctly by the models.
Baselines. We compare CWT with other input transformation-based methods. Specifically, the image scale-based methods (DIM [50], SIM [20]), the mixed image-based methods (Admix [42], US-MM [40]), and the block-based methods (MaskBlock [7], BSR [39]). For fairness, all the input transformations are integrated into MI-FGSM [4].
Evaluation Settings. We set the maximum perturbation $\epsilon=16/255$, the number of iterations $epoch=10$, the step size $\alpha=\epsilon/epoch$, the batch size $batchsize=8$, and the decay factor $\mu=1$ for MI-FGSM. For our method, CWT generates 20 scaled copies per iteration, divides the image into 2x2 blocks, applies a scaling factor ranging from 1.0 to 1.3, and applies a maximum rotation angle of $26^{\circ}$, selectively rotating $k=2$ blocks. For other methods, we follow the parameters reported in the original papers.
@@ 4.2 Evaluations on CNN-based Models
We first evaluate the transferability of adversarial examples generated by various attacks with input transformations when the surrogate model is based on a CNN architecture. Specifically, we generate adversarial examples using RN-18, RN-101, RX-50, and DN-121 as surrogate models and evaluate their attack success rates (ASR) across eight target models. The results are summarized in Table 1, where the first column indicates the surrogate model, the second column lists the attack methods, and the remaining columns present the ASR of different classification models under attack and the last two columns are the mean ASR and standard deviation, respectively.
Our proposed method consistently achieves state-of-the-art performance, with the highest mean ASR and lowest standard deviation across almost all experimental settings. Notably, when the surrogate model is RN-18, our method achieves the highest ASR across all target models. Furthermore, when the target models are ViT-B and PiT-B, our method surpasses the previous state-of-the-art BSR more than 5% and outperforms other baseline methods by at least 17.4%. These results confirm the effectiveness of our approach in generating transferable adversarial examples when the surrogate model is CNN-based.
@@ 4.3 Evaluations on Transformer-based models
Model Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$) Mean ($\uparrow$) Std. Dev. ($\downarrow$)
ViT-B DIM 56.9 47.8 49.6 59.7 89.5* 54.0 55.0 61.7 59.3 13.1
@@
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
@@
Attack RN-18 ($\uparrow$) RN-101 ($\uparrow$) RX-50 ($\uparrow$) DN-121 ($\uparrow$) ViT-B ($\uparrow$) PiT-B ($\uparrow$) ViF-S ($\uparrow$) Swin-T ($\uparrow$)
@@
CWT 40.5 56.7 38.4 41.3 43.2 43.2 48.7 37.8
Table 4: Attack success rates (%) of adversarial examples generated using various attack methods across eight classification models under Diffure. The classifier is RN-101.
Building on the results observed with CNN-based models, we next evaluate our method’s performance when the surrogate model is based on a transformer architecture. Specifically, we use Transformer-based models, including ViT-B, PiT-B, ViF-S, and Swin-T, as surrogate models to generate adversarial examples and evaluate their transferability. The results are summarized in Table 2.
From the results, we observe that our method also achieves SOTA performance in terms of both mean ASR and standard deviation across most of experimental settings. Notably, when the surrogate model is ViT-B, our method does not achieve the highest ASR on the original classification model (ViT-B itself) but consistently achieves the best performance on all other classification models. This demonstrates the robustness and generalizability of our approach, as it effectively transfers adversarial examples to different target architectures, even when the surrogate model has inherent challenges in transferability.